OpenStudy (anonymous):

solving with substitution x+y=11 -x=-y-9

OpenStudy (anonymous):

easier if you write the second equation as \[x=y+9\] then substitute in the first to get \[y+9+y=11\] \[2x+9=11\] \[2y=2\] \[y=1\]

OpenStudy (anonymous):

x+y=11 -x=-y-9 --> x = y+9 y+9+y=11 2y=2 y=1 x=10

OpenStudy (lalaly):

x+y=11....(1) -x=-y-9....(2) from (2) you get x=y+9 in (1) y+9+y=11 2y=2 y=1 x=1+9 x=10

OpenStudy (anonymous):

satellite meant \[2y+9=11\] on his second line. But the answer remains the same :)